# FEALPy v3: A Cross-platform Intelligent Numerical Simulation Engine

### v1

@@ 1 Introduction
Mature numerical PDE software such as iFEM[Chen:2008ifem], AFEPack[Cai:2024afepack], MFEM[Anderson:2021mfem], FEniCSx[BarattaEtal2023, ScroggsEtal2022, BasixJoss, AlnaesEtal2014], deal.II[Arndt:2019dealII], PHG, and OpenFOAM[Jasak:2009OpenFOAM] offer powerful and well-established infrastructures for numerical simulation. These software systems have greatly advanced the development and deployment of high-performance PDE solvers across many scientific and engineering domains. While emerging demands—such as unified tensor representations, interaction with modern DL toolchains, and cross-backend portability—call for complementary approaches that extend beyond the traditional designs of existing packages.
Motivated by this observation, we have developed FEALPy, a tensor-centric numerical simulation engine designed for flexibility, interoperability, and extensibility. FEALPy builds a unified abstraction layer over multiple tensor computation backends (including NumPy[HARRIS2020-NUMPY], PyTorch[PASZKE2019-PYTORCH], JAX[JAX2018] and more in the future), enabling consistent data flow across heterogeneous hardware architectures. Its modular design decouples low-level tensor operations from higher-level algorithmic modules—such as meshes, function spaces, discretization schemes, and solvers—allowing users to implement complex simulation workflows with concise and reusable code. By adopting a unified tensor interface, FEALPy provides a common computational foundation for both classical numerical algorithms and differentiable programming. It supports automatic differentiation (AD) throughout the entire pipeline, making it possible to embed finite element solvers as differentiable layers within NNs. This integration bridges the gap between traditional physics-based modeling and data-driven learning, advancing the convergence of numerical simulation and modern AI.
@@ 2.1 Overview
FEALPy is a modular numerical simulation engine built around tensorized computing. Its architecture separates mathematical abstractions from computational details so that numerical algorithms remain independent of specific hardware or software backends. Based on the dependencies between tensors, geometry, algorithms and applications, from bottom to top, we divide FEALPy into four layers: tensor layer, commonality layer, algorithm layer and field layer, see Fig. 1. The tensor layer provides an abstract interface to meet the needs of all upper modules that are based on tensor computations; the commonality layer provides basic data structures and algorithms, such as meshes, function spaces, sampling, integration and solvers; the algorithm layer provides many classic numerical algorithm frameworks, such as FEM, FDM, FVM, SPH; the field layer enters each specific field and provides common material settings, physical models, computational models and examples.
@@ 2.2 Backend Module and Hardware Adaptation
The backend layer is the foundation of FEALPy’s interoperability. It provides a backend manager (BM) that standardizes tensor operation interfaces across multiple frameworks. Every tensor operation in FEALPy—such as array creation, broadcasting, reduction, or Einstein summation—is routed through this manager, which dynamically dispatches calls to the active backend. This design mirrors the multi-backend philosophy adopted by TensorLy[TENSORLY], where the same high-level logic can execute on CPUs or GPUs without code modification.
Backend layer exposes a unified API compatible with the Python Array API Standard[ARRAYAPI2024]. It abstracts the complexities of managing different tensor computation backends, allowing FEALPy to integrate various libraries such as NumPy, PyTorch, JAX, MindSpore, PaddlePaddle, and others. This flexibility ensures that users can seamlessly adapt their simulations to the most suitable tensor backend based on specific hardware requirements. In addition, as all higher-level modules operate solely on BM to access APIs, FEALPy ensures that algorithms written once in terms of abstract tensor operations can be executed on any supported backend. This not only simplifies development by providing a unified API, but also ensures long-term maintainability and adaptability as new backends and hardware architectures emerge.
@@ 2.3 Common Modules with Unified Interfaces
  • Sparse: A variety of sparse formats, including CSR, CSC, and COO, supports multiple backends.
  • Quadrature: Generate quadrature rules for numerical integration.
  • Plotting: Supplies lightweight visualization tools for quick inspection of meshes and values attached to their entities.
@@ 3 Implementation Details and Programming Techniques
3.1 Backend Manager and Adaptors
A significant advantage of FEALPy’s Tensor Backend System is its seamless backend switching capability. Users can optimize simulations for different computing environments—whether they are running simulations on CPUs, GPUs, or specialized hardware like TPUs—without modifying their codebase. The Lst. 1 snippet demonstrates how FEALPy enables users to switch between different tensor backends with minimal effort. This seamless switching capability allows developers to experiment with various backends and select the one that offers the best performance for their hardware, ensuring optimized simulation results.
@@ 3.1 Backend Manager and Adaptors
Listing 1: Switching between different tensor backends: By replacing the backend name passed to the set_backend function, the object type of the output tensor changes. bm here refers to the backend manager, which provides standard array APIs.
How does the Backend module build a tensor abstraction layer so that users can operate multiple tensor libraries through almost the same interface? Constructed by a backend manager and several backend adaptors, the Backend module has two working steps: transmission and translation.
BackendManager is a router that calls tensor operations of the selected version, and the corresponding adaptor is initialized if it has not been loaded yet. This lazy loading feature avoids unnecessary importation.
While Adaptors are translators that simulate the standard interface based on the corresponding tensor library. Three cases are considered in each interface implementation: (1) satisfied by a function in the tensor library: then the function is directly referenced in the adapter; (2) satisfied at the functional level, but with a different function name: then the function is referenced in the adapter via a mapping table; (3) not satisfied: then manual implementation based on the corresponding tensor library is required in the adapter.
This super-polymorphic feature between different tensor libraries will give the same ability to switch between tensor libraries for almost all implementations of high-level modules, such as those presented in the following subsections.
@@ 3.2 Mesh Data Structure Based on Tensors
Based on the tensor abstraction interface, we designed the mesh data structure in FEALPy carefully. The Cartesian coordinate of points are recorded in a floating-point tensor, which is then combined into higher-dimensional shapes through a series of non-negative integer tensors. We adopt the notation shown in Tab. 1.
@@
Two vertices can form a line segment by connecting them in a directed way, so a 2-tuple containing the global indices of the start and end points can represent an edge. Edges of a mesh can be represented as a tensor with shape [NE, 2] in almost all unstructured mesh types in FEALPy. 2D triangles are constructed in FEALPy by connecting three vertices in counter-clockwise order, thus [num_triangles, 3] records a collection of them in a mesh. See Lst.2 as an example for a mesh containing 2 triangles divided from a square, and the rules for other shapes are similar.
@@
Polygons describe a group of shapes with inconsistent numbers of edges, and these shapes cannot be combined to form 2D tensors similar to the rules for triangles and quadrangles. All vertices from multiple polygons are directly arranged in a 1D tensor, and a pointer tensor is used to record the starting position (with an additional total number of vertices as the last element) of each polygon, as shown in Fig. 3. The length of this pointer tensor is exactly num_polygons + 1, like the indptr in CSR format sparse matrices.
@@
In the case of unstructured meshes, FEALPy always forms these topological 2-dimensional shapes by ordering the vertices in a circular manner, so it is natural to define the orientation by the ordering direction. The right-hand rule is used to define the orientation in FEALPy. See Fig. 3.
@@
A deduplication process is needed in the mesh construction, because edges and faces are bound to have duplicates if they are simply collected from all cells. In particular, a face is definitely adjacent to two cells wanted by FEALPy to construct a topological relationship named face_to_cell in the shape of [NF, 4]. Tab. 2 shows what this four columns mean. FEALPy deduplicates the lower-dimensional entities using lexsort, through which the first and the last occurrences of the same faces are extracted from the total. Since the right-hand rule is used for 2D face orientation, the cell where a face firstly appears is the cell the face faces to; and since the counter-clockwise rule is used for 1D edge orientation, the cell where an edge firstly appears is the cell on the left side of the edge.
@@
Table 2: Definition of the 4 columns in face_to_cell. Once the indices in total faces are extracted, cell indices and local indices can be deduced by floor division and modulo.
@@
By grouping various entities together and adding related methods, a mesh class can be constructed. All supported mesh types are listed in Tab.3. The Mesh module is designed to maximize the reuse of methods through an reasonable inheritance structure and keep the interface of methods consistent between different mesh types. Therefore, meshes can be simply replaced without the need to adapt the rest of the script. See Lst.3.
@@
Half Edge 1 Edge 1 Dart Mesh 2/3
Interval 1 Triangle 2 Uniform Mesh 1/2/3
Quadrangle 2 Polygon 2
Tetrahedron 3 Hexahedron 3
Prism 3 Node 0
@@
Table 3: Mesh types of unstructured (left and middle columns) and others (right column) in FEALPy and their supported TDs.
@@ 3.3 Basis and DoF Management
For most finite element methods or their variants, DoFs are arranged by interpolation points on the mesh entities, and there are two main types of DoF management provided in FEALPy in this way. The first type is continuous, where adjacent mesh cells may share a part of DoFs at their interface. Section 6 of [CHEN2023GEO] has discussed the global indexing rules for Lagrange interpolation points in detail, which is actually the main idea behind FEALPy’s implementation of various finite element DoF management. In order to provide a non-repeating numbering that is as consistent as possible for all mesh entities, we adopt the order of DoFs from low-dimensional entities to high-dimensional entities in the order of "node-edge-face-cell", as shown in Fig. 4. When numbering the DoFs inside two adjacent cells, the DoFs at their interface are already numbered. The second type is discontinuous, where adjacent mesh cells do not share any DoF. Non-repeating numbers can be generated directly from the correspondence between cells and their local DoFs.
@@
Based on different basis function implementations and mappings of local-global DoFs, FEALPy already supports a variety of function spaces summarized in Tab. 4.
@@
Bernstein Any [KIRBY2011-Bernstein]
$C^{m}$ Conforming 2, 3 [CHENLONG2021, HUJUN2023, CHEN2025]
Hu-Zhang 2, 3 [HU2015-HuZhang]
First Nédélec 2, 3 [NEDELEC1986]
Second Nédélec 2, 3 [NEDELEC1986], [CHEN2023GEO]
Raviart-Thomas 2, 3 [RAVIART1977]
Brezzi-Douglas-Marini 2, 3 [BREZZI1985-BDM], [CHEN2023GEO]
@@ 3.4 Sampling and Quadrature
Sampling and quadrature are also essential tools for many numerical methods. The Quadrature module stores precomputed barycentric coordinates of quadrature points and their weights generated from symmetric quadrature rules on simplexes [WILLIAMS201418], including line segments, triangles, and tetrahedrons. While quadrature points on $d$-hypercubes, such as quadrangles and hexahedrons, are generated via tensor products of 1D Gauss-Legendre quadrature rules. Similarly, triangular prisms and pyramids can also get their quadrature rules via tensor products. Lst.4 gives an example.
@@
Listing 4: Quadrature points on quadrangles can be generated via tensor products of two 1D Gauss-Legendre quadrature rules.
Samplers are designed for generating equidistant or random collocation points for function learning and operator learning. Sampling can be done mesh-independent, which is different from the Quadrature module, as a path of circumventing the curse of dimensionality.
@@ 3.5 Sparse Format and Matrix Assembly
Sparse COO and CSR formats are implemented in FEALPy and can work on any tensor library supported by FEALPy, as they are completely based on the tensor data structure and is decoupled with any specific backends.
@@
Let d be the dimension (number of axes) and nnz be the number of non-zeros. In both COO and CSR, FEALPy allows None data (matrices with bool elements, for sparse adjacency) and data with a batch axis as the 0-axis, as shown in Fig. 5.
@@
Combining the cell matrix from Integrator and the DoF management from the FunctionSpace, Forms in the FEM module manages to arrange the non-zero elements in sparse linear systems leveraging the Sparse module. Lst.5 gives an example to explain the arrangement.
@@
Listing 5: Assembly of global stiffness matrices. The local_shape equals to (num_cells, num_quadpts, num_basis, num_basis) which stores the shape of cell_matrix; ue2dof and ve2dof are DoF maps, in the shape of (num_cells, num_dofs).
Unfortunately, to our best knowledge, matrix-vector multiplication of batched sparse matrices is not well supported now by any tensor computation library at the C/C++ level. Although implementations based on standard tensor computation at the Python level can output correct results, a significant loss of efficiency can be observed in solvers.

### v2

# FEALPy: A Cross-platform Intelligent Numerical Simulation Engine

@@ 1 Introduction
Mature numerical PDE software such as iFEM[11], AFEPack[7], MFEM[2], FEniCSx[4, 33, 32, 1], deal.II[3], PHG, and OpenFOAM[20] offer powerful and well-established infrastructures for numerical simulation. These software systems have greatly advanced the development and deployment of high-performance PDE solvers across many scientific and engineering domains. While emerging demands—such as unified tensor representations, interaction with modern DL toolchains, and cross-backend portability—call for complementary approaches that extend beyond the traditional designs of existing packages.
Motivated by this observation, we have developed FEALPy, a tensor-centric numerical simulation engine designed for flexibility, interoperability, and extensibility. FEALPy builds a unified abstraction layer over multiple tensor computation backends (including NumPy[15], PyTorch[28], JAX[5] and more in the future), enabling consistent data flow across heterogeneous hardware architectures. Its modular design decouples low-level tensor operations from higher-level algorithmic modules—such as meshes, function spaces, discretization schemes, and solvers—allowing users to implement complex simulation workflows with concise and reusable code. By adopting a unified tensor interface, FEALPy provides a common computational foundation for both classical numerical algorithms and differentiable programming. It supports automatic differentiation (AD) throughout the entire pipeline, making it possible to embed finite element solvers as differentiable layers within NNs. This integration bridges the gap between traditional physics-based modeling and data-driven learning, advancing the convergence of numerical simulation and modern AI.
FEALPy is fully open-source under the GNU GPL v3.0 license, with code hosted on GitHub at https://github.com/weihuayi/fealpy.
@@ 2.1 Overview
FEALPy is a modular numerical simulation engine built around tensorized computing. Its architecture separates mathematical abstractions from computational details so that numerical algorithms remain independent of specific hardware or software backends. Based on the dependencies between tensors, geometry, algorithms and applications, from bottom to top, we divide FEALPy into four layers: tensor layer, commonality layer, algorithm layer and field layer, see Figure 1. The tensor layer provides an abstract interface to meet the needs of all upper modules that are based on tensor computations; the commonality layer provides basic data structures and algorithms, such as meshes, function spaces, sampling, integration and solvers; the algorithm layer provides many classic numerical algorithm frameworks, such as FEM, FDM, FVM, SPH; the field layer enters each specific field and provides common material settings, physical models, computational models and examples.
@@ 2.2 Backend Module and Hardware Adaptation
The backend layer is the foundation of FEALPy’s interoperability. It provides a backend manager (BM) that standardizes tensor operation interfaces across multiple frameworks. Every tensor operation in FEALPy—such as array creation, broadcasting, reduction, or Einstein summation—is routed through this manager, which dynamically dispatches calls to the active backend. This design mirrors the multi-backend philosophy adopted by TensorLy[22], where the same high-level logic can execute on CPUs or GPUs without code modification.
Backend layer exposes a unified API compatible with the Python Array API Standard[29]. It abstracts the complexities of managing different tensor computation backends, allowing FEALPy to integrate various libraries such as NumPy, PyTorch, JAX, MindSpore, PaddlePaddle, and others. This flexibility ensures that users can seamlessly adapt their simulations to the most suitable tensor backend based on specific hardware requirements. In addition, as all higher-level modules operate solely on BM to access APIs, FEALPy ensures that algorithms written once in terms of abstract tensor operations can be executed on any supported backend. This not only simplifies development by providing a unified API, but also ensures long-term maintainability and adaptability as new backends and hardware architectures emerge.
@@ 2.3 Common Modules with Unified Interfaces
  • Sparse: A variety of sparse formats (CSR, CSC, and COO) supporting multiple backends.
  • Quadrature: Quadrature rules for numerical integration.
  • Plotting: Lightweight visualization tools for quick inspection of meshes and values attached to their entities.
@@ 3 Implementation Details and Programming Techniques
3.1 Backend Manager and Adapters
A significant advantage of FEALPy’s Tensor Backend System is its seamless backend switching capability. Users can optimize simulations for different computing environments—whether they are running simulations on CPUs, GPUs, or specialized hardware like TPUs—without modifying their codebase. Listing 1 snippet demonstrates how FEALPy enables users to switch between different tensor backends with minimal effort. This seamless switching capability allows developers to experiment with various backends and select the one that offers the best performance for their hardware, ensuring optimized simulation results.
@@ 3.1 Backend Manager and Adapters
Listing 1: Switching between different tensor backends: By replacing the backend name passed to the
set_backend function, the object type of the output tensor changes. bm here refers to the backend manager, which provides standard array APIs.
The Backend module provides a unified tensor abstraction layer that enables users to operate different tensor libraries (e.g., NumPy, PyTorch) through a nearly identical interface. It consists of two main components: a BackendManager and a set of backend adapters. The workflow involves two stages: transmission and translation.
The BackendManager acts as a dispatcher. When a tensor operation is requested, it routes the call to the currently selected backend. If the corresponding adapter has not yet been initialized, it is loaded at that moment. This lazy-loading mechanism prevents unnecessary imports and reduces initialization overhead.
Each adapter implements the Python Array API Standard for a specific tensor library. For every interface function required by the standard, the adapter determines whether and how the target tensor library provides equivalent functionality. In this context, “satisfied” means that the tensor library already provides an operation that matches the required functionality of the standard, either directly or with minor adaptation. “Not satisfied” means that the tensor library does not provide such functionality, and the operation must therefore be implemented manually within the adapter.
Specifically, three situations are considered for each interface function:
  • Directly satisfied: The tensor library provides a function whose name and behavior are consistent with the required interface. In this case, the adapter directly references this function.
  • Functionally satisfied but with different naming or signature: The tensor library provides equivalent functionality, but the function name or argument format differs from the standard interface. In this case, the adapter maps the standard interface function to the corresponding library function using a predefined mapping rule.
  • Not satisfied: The tensor library does not provide a function that fulfills the required functionality. In this case, the adapter implements the operation explicitly using available primitives from the tensor library.
Through this mechanism, FEALPy ensures that high-level modules remain consistent across different backends, while backend-specific differences are encapsulated within the adapters.
@@ 3.2 Mesh Data Structure Based on Tensors
Based on the tensor abstraction interface, we designed the mesh data structure in FEALPy carefully. The Cartesian coordinate of points are recorded in a floating-point tensor, which is then combined into higher-dimensional shapes through a series of non-negative integer tensors. We adopt the notation shown in Table 1.
@@
Two vertices can form a line segment by connecting them in a directed way, so a 2-tuple containing the global indices of the start and end points can represent an edge. Edges of a mesh can be represented as a tensor with shape [NE, 2] in almost all unstructured mesh types in FEALPy. 2D triangles are constructed in FEALPy by connecting three vertices in counter-clockwise order, thus [num_triangles, 3] records a collection of them in a mesh. Listing 2 is an example for a mesh containing 2 triangles divided from a square, and the rules for other shapes are similar.
@@
Polygons describe a group of shapes with inconsistent numbers of edges, and these shapes cannot be combined to form 2D tensors similar to the rules for triangles and quadrangles. All vertices from multiple polygons are directly arranged in a 1D tensor, and a pointer tensor is used to record the starting position (with an additional total number of vertices as the last element) of each polygon, as shown in Figure  3. The length of this pointer tensor is exactly num_polygons + 1, like the indptr in CSR format sparse matrices.
@@
In the case of unstructured meshes, FEALPy always forms these topological 2-dimensional shapes by ordering the vertices in a circular manner, so it is natural to define the orientation by the ordering direction. The right-hand rule is used to define the orientation in FEALPy. See Figure 3.
@@
Some topological relationships between mesh entities are available in FEALPy, such as face_to_cell in the shape of [NF, 4]. A deduplication process is needed for constructing these relationships, because lower-dimensional entities, e.g. edges and faces, are bound to have duplicates if they are simply collected from all higher-dimensional cells. Table 2 shows what the four columns of face_to_cell mean. FEALPy deduplicates the lower-dimensional entities using lexsort, through which the first and the last occurrences of the same faces are extracted from the total. Since the right-hand rule is used for 2D face orientation, the cell where a face firstly appears is the cell the face faces to; and since the counter-clockwise rule is used for 1D edge orientation, the cell where an edge firstly appears is the cell on the left side of the edge.
@@
Table 2: Definition of the 4 columns in face_to_cell. Each row corresponds to a face after deduplication, and the first two columns represent the indices of the two cells that this face belongs to, while the last two columns represent the local indices of this face in these two cells.
@@
By grouping various entities together and adding related methods, a mesh class can be constructed. All supported mesh types are listed in Table 3. The Mesh module is designed to maximize the reuse of methods through an reasonable inheritance structure and keep the interface of methods consistent between different mesh types. Therefore, meshes can be simply replaced without the need to adapt the rest of the script.
Listing 3 shows mesh types that currently supported in FEALPy. And we would like to make some extra explanations here:
  • The NodeMesh is a special mesh type that only contains nodes without any higher-dimensional entities, and the topological relationships between nodes are dynamically constructed, for example, by connecting adjacent nodes leveraging a KD-tree. This is particularly useful for methods like smooth particle hydrodynamics (SPH) and deep learning methods that operate on point clouds.
  • Each edge of a mesh can be divided into a pair of opposite edges if the directionality is considered, where each in the pair is called a ’half edge’. Therefore, the entire mesh, as a HalfEdgeMesh, has a total of $2\texttt{NE}$ half edges.
@@
Interval 1 Edge 1 Node Mesh 0
Quadrangle 2 Triangle 2 HalfEdge Mesh 1
Tetrahedron 3 Polygon 2 Dart Mesh 2/3
Prism 3 Hexahedron 3 Uniform Mesh 1/2/3
@@
Table 3: Types of unstructured meshes (left and middle columns) and others (right column) in FEALPy and their supported TDs.
@@ 3.3 Basis and DoF Management
For most finite element methods or their variants, DoFs are arranged by interpolation points on the mesh entities, and there are two main types of DoF management provided in FEALPy in this way. The first type is continuous, where adjacent mesh cells may share a part of DoFs at their interface. Section 6 of [9] has discussed the global indexing rules for Lagrange interpolation points in detail, which is actually the main idea behind FEALPy’s implementation of various finite element DoF management. In order to provide a non-repeating numbering that is as consistent as possible for all mesh entities, we adopt the order of DoFs from low-dimensional entities to high-dimensional entities in the order of "node-edge-face-cell", as shown in Figure 4. When numbering the DoFs inside two adjacent cells, the DoFs at their interface are already numbered. The second type is discontinuous, where adjacent mesh cells do not share any DoF. Non-repeating numbers can be generated directly from the correspondence between cells and their local DoFs.
@@
Based on different basis function implementations and mappings of local-global DoFs, FEALPy already supports a variety of function spaces summarized in Table  4.
@@
Bernstein Any [21]
$C^{m}$ Conforming 2, 3 [10, 17, 8]
Hu-Zhang 2, 3 [18]
First Nédélec 2, 3 [27]
Second Nédélec 2, 3 [27], [9]
Raviart-Thomas 2, 3 [30]
Brezzi-Douglas-Marini 2, 3 [6], [9]
@@
Tensor function spaces.
Given a scalar function space, the TensorFunctionSpace class in FEALPy constructs a tensor-valued function space by assigning a prescribed tensor shape to each scalar degree of freedom (DoF). A leading or trailing ‘-1‘ in the shape specifies the ordering of the tensor DoFs:
  • A leading ‘-1‘ indicates that the scalar DoF index serves as the outer index, so that the tensor component index varies fastest.
  • A trailing ‘-1‘ indicates that the scalar DoF index serves as the inner index, so that the scalar DoF index varies fastest within each tensor component.
Listing 4 illustrates the construction of a tensor function space with shape (-1, 2)—a 2-component vector-valued function space where the scalar DoF index is the outer index.
1from fealpy.functionspace import (
2 LagrangeFESpace, TensorFunctionSpace
3)
4...
5scalar_space = LagrangeFESpace(mesh=mesh, p=1)
6tensor_space = TensorFunctionSpace(scalar_space, shape=(-1, 2))
Listing 4: A tensor-valued function space was constructed by shape (-1, 2), from a given scalar space. The leading -1 means that the scalar DoF index serves as the outer index, so that the tensor component index varies fastest.
@@ 3.4 Sampling and Quadrature
Sampling and quadrature are also essential tools for many numerical methods. The Quadrature module stores precomputed barycentric coordinates of quadrature points and their weights generated from symmetric quadrature rules on simplexes [35], including line segments, triangles, and tetrahedrons. While quadrature points on $d$-hypercubes, such as quadrangles and hexahedrons, are generated via tensor products of 1D Gauss-Legendre quadrature rules (Listing 5). Similarly, triangular prisms and pyramids can also get their quadrature rules via tensor products.
@@
Listing 5: Quadrature points on quadrangles can be generated via tensor products of two 1D Gauss-Legendre quadrature rules.
Samplers are designed for generating equidistant or random collocation points for deep learning methods. For instance, ISampler and BoxBoundarySampler can generate sampling points in the interior and on the boundary of a given rectangular region respectively; while MeshSampler can draw sampling points on given homogeneous mesh entities.
@@ 3.5 Sparse Format and Matrix Assembly
FEALPy implements sparse matrix representations in COO and CSR formats. These implementations are backend-agnostic: all sparse data structures are constructed entirely from the tensor abstraction layer provided by the Backend module, and therefore operate uniformly across all supported tensor libraries.
@@
Let $d$ denote the number of tensor axes and $\mathrm{nnz}$ the number of nonzero entries. In addition to standard COO and CSR definitions, FEALPy extends these formats in two aspects.
First, the data field may be None, which corresponds to structure-only sparse matrices (e.g., Boolean adjacency matrices), where only index information is required.
Second, the data tensor may contain a leading batch axis. Specifically, if the sparse tensor represents a batch of matrices sharing the same sparsity pattern, the index tensor has shape $(d,\mathrm{nnz})$, while the data tensor has shape $(b,\mathrm{nnz})$, where $b$ is the batch size. This hybrid representation (Figure 5) enables efficient storage of multiple sparse matrices with identical nonzero structure but different numerical values. Such a design is particularly useful in finite element computations where multiple linear systems with varying coefficients share the same connectivity pattern.
@@
In the finite element module, global sparse matrices are assembled by combining local element matrices produced by the Integrator with degree-of-freedom (DoF) mappings provided by the FunctionSpace. The Forms component coordinates this process using the Sparse module.
Let cell_matrix denote the local stiffness matrices with shape
$(\texttt{num\_cells},~\texttt{num\_quadpts},~\texttt{num\_basis},~\texttt{num\_basis}),$
and let ue2dof and ve2dof be the DoF mappings from local basis functions to global indices, both of shape $(\texttt{num\_cells},~\texttt{num\_dofs}).$ The global row and column indices of all nonzero entries are generated by broadcasting the DoF mappings to match the local tensor shape. Listing 6 illustrates this procedure.
@@
Listing 6: Assembly of global stiffness matrices. The local_shape equals to (num_cells, num_quadpts, num_basis, num_basis) which stores the shape of cell_matrix; ue2dof and ve2dof are DoF maps, in the shape of (num_cells, num_dofs).
The broadcasted tensors I and J represent the global row and column indices of each local contribution. After flattening, these indices are stacked to form the COO index tensor of shape $(2,\mathrm{nnz})$, where $\mathrm{nnz}$ equals the total number of local contributions. The local matrices are reshaped accordingly to produce the associated COO data tensor. The global sparse matrix is then constructed by accumulating these contributions via the add operation.
For batched problems, the local contribution tensor acquires an additional leading dimension corresponding to the batch size, resulting in a sparse tensor with shape $(b,\mathrm{nnz})$ in the data field.
Current limitations.
At present, mainstream tensor computation libraries provide limited low-level (C/C++) support for batched sparse matrix–vector multiplication when matrices share identical sparsity patterns. Although correct implementations can be constructed at the Python level using standard tensor operations, these approaches typically rely on repeated sparse operations or indirect indexing, which introduces additional overhead. As a result, solver performance may degrade in large-scale batched computations.
Improving backend-level support for batched sparse linear algebra remains an important direction for future development. And note that our unified tensor abstraction layer allows us to integrate future improvements or user C/C++ extensions (by updating the existing adapters or adding a new backend adapter) without changing the high-level algorithm code. So the current framework will not prevent future improvements in this area.